OpenStudy (anonymous):

Find each product. Simplify if possible. (-7/8)(-4/5)

OpenStudy (anonymous):

0.7

OpenStudy (idealist10):

(-7/8)(-4/5)=28/40=7/10 Answer: 7/10 Remember, negative times negative gives you positive.

OpenStudy (anonymous):

can yu explain

OpenStudy (idealist10):

Well, you just multiply numerator with numerator and multiply denominator with denominator. Whatever you get, you simplify.